calculate the molality,molarity, and mole fraction of NH3 which is 8% by mass aq. solution (d=0.9651g/ml
Answer #1

Assuming 1 L of solution

Mass of solution = density * volume = 0.9651 g/ml * 1 L * 1000 ml/1 L = 965.1 g

Mass of NH3 = 8% of 965.1 g = 8/100 * 965.1 = 77.21 g

Molar mass of NH3 = Molar mass of N + 3 * Molar mass of H = 14 + 3 * 1 = 17 g/mol

Number of moles of NH3 = Mass/Molar mass = 77.21/17 = 4.54 moles

Molarity = Number of moles of solute/Volume of solution (in L) = 4.54 moles/1 L = 4.54 M

Mass of solvent = Mass of solution - Mass of NH3 = 965.1 - 77.2 = 887.89 grams

Molality = Number of moles of solute/Mass of solvent in (kg) = 4.54/0.88789 = 5.113 m

Number of moles of water = Mass/molar mass = 887.89/18 = 49.32 moles

mole-fraction of NH3 = (number of moles of NH3)/(number of moles of NH3 + number of moles of water) = 4.54/(4.54+49.32) = 0.0843

Note - Post any doubts/queries in comments section.